Number of solution(s) of the equation           |z|^2 + 7z = 0 is/are

Grade:11

2 Answers

Deepak Kumar Shringi
5 years ago
put z=x+iy
and |z|=(x2+y2)1/2
then real part and img part to both side equate
Aditya Gupta
5 years ago
We see that
-|z|^2/7= z, which means that z is real and less than or equal to zero. So modz= -z
So equation becomes
-z^2=7z
Or z= 0 or -7